# Weighing the Axion with Muon Haloscopy

### 06 Dec 2023 v35

@@ Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the Standard Model of particle physics. We argue that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument yields a percent level prediction for the mass of the axion provided that dark matter is made of axions. An analysis of charge asymmetry in kaon decays suggests that at least in the local halo dark matter is made of axions and that axions from the local halo are responsible for the observed violation by these reactions of the combined charge conjugation and spatial inversion symmetry operation. Tabletop experiments to directly detect dark matter in the form of axions with the predicted mass are proposed.
Introduction—The spin of the muon precesses in a magnetic field $B$ at frequency $\omega_{a}=a_{\mu}(e/m_{\mu})B$, where, $a_{\mu}\approx\alpha/(2\pi)\approx 1.16\times 10^{-3}$ is the electromagnetic leading order contribution to spin precession, and $e/m_{\mu}=2\pi\times 135~{}{\rm MHz/T}$ is the charge to mass ratio of the muon [1, 2, 3, 4, 5]. Compared with the electron, the muon is $(m_{\mu}/m_{e})^{2}\approx 43,000$ times more sensitive to the hadronic leading order contribution $a_{\mu}^{{\rm HLO}}\approx 700\times 10^{-10}$ to spin precession [5, 6]. In this Article, we show that axions in the local dark matter halo of the galaxy shift $a_{\mu}$ and $a_{\mu}^{{\rm HLO}}$ from their standard model values. Further, we argue that these axion-induced shifts resolve the tension between measurements of the muon spin precession frequency [2, 3, 4] and calculations of the frequency within the Standard Model [6, 7, 8, 9, 10, 11] provided that dark matter in the local halo is made of axions as suggested by analysis of charge asymmetry between the kaon decay $K^{0}_{L}\rightarrow\pi^{+}e^{-}\overline{\nu}_{e}$ and its charge-parity conjugate $K^{0}_{L}\rightarrow\pi^{-}e^{+}\nu_{e}$ [12, 13, 14, 15, 16]. Finally, we give a physical picture for the nature of dark matter that leads to a prediction for the mass of the axion with percent level precision and we sketch table-top experiments for testing this prediction.
@@
Figure 1: Halo axion $A$ collides with virtual photon or vector meson dressing vertex of positive muon $\mu^{+}$ and magnetic field $B$.
In contrast to the local axion energy density, the axion mass $m_{A}$ and decay constant $f_{A}$ only enter the kaon charge asymmetry through the product $m_{A}f_{A}=\sqrt{\chi_{\rm QCD}}=(5.69\pm 0.05)\times 10^{-3}~{}{\rm GeV^{2}}$ which is fixed by the topological susceptibility $\chi_{\rm QCD}$ of the quantum chromodynamic vacuum [27]. Using the kaon parameters $L_{K}=500~{}{\rm cm}$ and $A_{L}(e)=(3.32\pm 0.08)\times 10^{-3}$ from the current best charge asymmetry measurement [17], the local axion energy density $\rho_{A}=(0.28\pm 0.03)~{}{\rm GeV/cm^{3}}$ that we find from Eq. (3) agrees with the best available estimate $\rho_{\rm DM}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ of the energy density of the local dark matter halo [28]. This agreement suggests that axions saturate the local dark matter halo and that dark matter from the local halo is the source of the charge asymmetry observed in kaon decays [29].
@@
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [35] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [36]. Using the muon parameters, we estimate the axion shift [37]
$\Delta a_{\mu}=(165\pm 10)\times 10^{-11}.$ (8)
The pleasing agreement of the axion shift $\Delta a_{\mu}$ and the observed gap $a_{\mu}({\rm Exp})-a_{\mu}({\rm SM})$ resolves the $3.4~{}\sigma$ tension between the Standard Model and the muon spin precession measurements.
Hadronic Haloscopy—Figure 1 also shows how axions from the local dark matter halo of the galaxy shift measurements of the leading order hadronic contribution $a_{\mu}^{\rm HLO}$ to muon spin precession [5, 6]. Acting as a background field, the axions couple to the $\rho^{0}$ vector meson that dominates the hadronic production from electron-positron annihilation whose cross-section enters the standard model calculation of $a_{\mu}^{\rm HLO}$ [38]. The axion coupling converts the short-lived $\rho^{0}$ meson, which produces mainly charged pion $\pi^{+}\pi^{-}$ pairs, into the longer lived and slighly heavier $\omega$ meson which prefers to produce a neutral pion $\pi^{0}$ in addition to the charged pion pair [39].
Dark matter-driven conversions of the neutral $\rho^{0}$ vector meson into the isoscalar $\omega$ vector meson reduce the resonant charged pion pair production cross-section that enters the resonant part $a_{\mu}^{\pi\pi}\approx 380\times 10^{-10}$ [8] of the hadronic leading order contribution $a_{\mu}^{\rm HLO}$ [40]. The axion-meson coupling $g_{A\rho\omega}=898~{}g_{A\gamma\gamma}=(6.11\pm 0.18)\times 10^{-8}~{}{\rm GeV}^{-1}$ [41] then leads to the shift [42]:
$\frac{\Delta a_{\mu}^{\pi\pi}}{a_{\mu}^{\pi\pi}}$ $=$ $m_{A}E_{\rho}g_{A\rho\omega}\tilde{\phi}_{A}(q_{A})$
$=$ $\frac{g_{A\rho\omega}}{g_{A\gamma\gamma}}\frac{\Delta a_{\mu}}{a_{\mu}}\frac{E_{\rho}}{m_{A}}\sqrt{\frac{1}{N_{\mu}}\frac{V_{\rho}}{V_{\mu}}\frac{T_{\rho}}{T_{\mu}}},$
where, $E_{\rho}$ is the vector meson energy in the axion rest frame, $V_{\rho}$ is the luminous volume created by the colliding electron and positron bunches, and $T_{\rho}$ is the time it takes the bunches to cross during the collision. The beam parameters give the axion shift [41, 43]
$\Delta a_{\mu}^{\pi\pi}$ $=$ $(19.6\pm 1.2)\times 10^{-11}\left(\frac{E_{\rho}}{1000~{}{\rm MeV}}\right)\left(\frac{\sigma_{x}}{100~{}{\rm\mu m}}\right)^{1/2}$ (10)
$\times$ $\left(\frac{\sigma_{y}}{10~{}{\rm\mu m}}\right)^{1/2}\left(\frac{\sigma_{z}}{1~{}{\rm cm}}\right),$
where, $\sigma_{x}$ is the width of the electron and positron bunches, $\sigma_{y}$ is their height, and $\sigma_{z}$ their length.
[…]
$=$ $(188\pm 37)\times 10^{-11}.$
Using Eq. (10) with the KLOE and CMD beam parameters gives a relative axion shift of roughly the right size to resolve the discrepancy [49]
[…]
$=$ $(167\pm 10)\times 10^{-11}.$
[…]
Proton Haloscopy— Inverting the process shown in Figure 1 about a horizontal axis and replacing the muon with a proton, we find the way that axions from the local dark matter halo of the galaxy shift the spin precession of protons in the molecules of a liquid during nuclear magnetic resonance [50]:
$\Delta g_{p}^{\prime}=\frac{\alpha}{\pi\epsilon}m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})=2\frac{\Delta a_{\mu}}{\epsilon}\sqrt{\frac{N_{p}}{N_{\mu}}\frac{V_{p}}{V_{\mu}}\frac{T_{p}}{T_{\mu}}},$ (13)
where, $\epsilon$ is the static dielectric constant of the liquid and $g_{p}^{\prime}\approx 5.58$ gives the proton spin precession frequency $\omega_{p}^{\prime}=(g_{p}^{\prime}/2)(e/m_{p})B\approx 2\pi\times 42.8~{}{\rm MHz/T}\times B$ for magnetic field $B$ [51]. Here, $N_{p}$ is the effective number of protons, $V_{p}$ is the volume containing the protons, and $T_{p}=1/\Delta\omega_{p}^{\prime}$ is the effective coherence time of the proton spin precession due to the broadening $\Delta\omega_{p}^{\prime}$ of the resonance by variations in the field strength over the sample volume. Straightforward estimates of the proton parameters give the slope of the axion shift of the proton spin precession frequency [52]
$\frac{1}{g_{p}^{\prime}}\frac{dg_{p}^{\prime}}{dT}$ $=$ $-(10.0\pm 0.6)~{}\frac{{\rm ppb}}{{\rm K}}\left(\frac{\epsilon}{78.4}\right)^{-1}\left(\frac{\rho}{1~{}{\rm g/cm^{3}}}\right)^{1/2}$ (14)
$\times$ $\left(\frac{\mu}{9~{}{\rm g/{mol~{}H}}}\right)^{-1/2}\left(\frac{T}{300~{}{\rm K}}\right)^{-3/2}\left(\frac{V_{p}}{{\rm 0.62~{}cm^{3}}}\right)$
$\times$ $\left(\frac{\Delta\omega_{p}^{\prime}}{2.6\times 10^{-7}~{}\omega_{p}^{\prime}}\right)^{-1/2},$
where, $T$ is the temperature, $\rho$ is the density, and $\mu$ is the mass of the liquid containing a mole of protons. Measurements of proton spin precession in liquids with the necessary precision to see the dark matter shift were achieved half a century ago [51]. Verifying the proton parameter dependence shown in Eq. (14) of the slope $dg_{p}^{\prime}/dT$ would directly confirm in a tabletop experiment the proposed mechanism for the shift of spin precession by axions from the local dark matter halo with the predicted mass.
Figure 2: Dark matter $A$ from the local halo of the galaxy converts infrared magnetic field $B_{\rm IR}$ into radio-frequency electric field $E_{\rm RF}$ that drives electron cyclotron motion and microwave cavity resonance.
Atomic Haloscopy—Replacing the positive muon in Figure 1 with the valence electron of an atom with large atomic number trapped in an optical lattice and replacing the external magnetic field with the internal magnetic field within the atom [53], we find the process that shifts the electromagnetic fine-structure coupling constant $\alpha$ felt by the electron during transitions among valence states:
$\frac{\Delta\alpha}{\alpha}=m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})=\frac{\Delta a_{\mu}}{a_{\mu}}\sqrt{\frac{N_{a}}{N_{\mu}}\frac{V_{a}}{V_{\mu}}\frac{T_{a}}{T_{\mu}}},$ (15)
where, $N_{a}$ is the number of atoms trapped in the optical lattice, $V_{a}$ is the trap volume, and $T_{a}$ is the coherence time for the motion of the ultracold atoms. The shift in $\alpha$ leads to a density-shift of the optical lattice clock frequency [54]:
$\frac{\Delta\nu}{\nu}$ $=$ $k_{\alpha}\frac{\Delta\alpha}{\alpha}=(11.8\pm 0.7)\times 10^{-17}\left(\frac{N_{a}}{100}\right)^{1/2}$
$\times$ $\left(\frac{V_{0}}{200~{}E_{R}}\right)^{-3/4}\left(\frac{T}{9~{}{\rm\mu K}}\right)^{1/4}\left(\frac{k_{\alpha}}{0.24}\right)$
$\times$ $\left(\frac{E_{R}}{h\times 2.0~{}{\rm kHz}}\right)^{-3/2}\left(\frac{m}{171~{}m_{p}}\right)^{-3/4}\left(\frac{w}{26.4~{}\lambda_{L}}\right)$
where, $k_{\alpha}$ is the sensitivity of the transition frequency to shifts in the fine-structure constant, $E_{R}=h^{2}/(2m\lambda_{L}^{2})$ is the kinetic energy with which the atoms of mass $m$ recoil when they spontaneously emit photons at the wavelength $\lambda_{L}$ of the optical lattice, $V_{0}$ is the lattice depth, $w$ is the lattice laser beam waist, and $T$ is the temperature of the atoms in the trap. By modulating the number of atoms in the trap, a density shift of the optical lattice clock frequency of this size can be readily observed [55].
Cyclotron Haloscopy—Figure 2 shows how dark matter drives electron cyclotron motion by converting infrared laser photons near the Compton wavelength $\lambda_{A}=2\pi/m_{A}=(2441\pm 19)~{}{\rm nm}$ into radio-frequency electromagnetic waves within a magnetically confined plasma. When the beat note between the infrared light and the Compton wavelength of dark matter hits the electron cyclotron resonance frequency $\omega_{c}=(e/m_{e})B=2\pi\times 28~{}{\rm GHz}~{}(B/1~{}{\rm T})$ for the magnetic field of strength $B$ that confines the plasma, the radio-frequency electric field created by the interference heats the electrons at the rate [56]
$P_{\rm RF}$ $=$ $\frac{P_{\rm IR}}{2}m_{A}^{4}g_{A\gamma\gamma}^{2}|\tilde{\phi}_{A}(q_{A})|^{2}$ (17)
$=$ $\frac{P_{\rm IR}}{2}\left(\frac{\Delta a_{\mu}}{a_{\mu}}\right)^{2}\frac{n_{e}V_{e}}{N_{\mu}}\frac{V_{e}}{V_{\mu}}\frac{T_{e}}{T_{\mu}}$
where, $P_{\rm IR}$ is the infrared power sent into the plasma, $n_{e}$ is the density of plasma electrons, $V_{e}$ is the volume of the infrared beam within the plasma, and $T_{e}$ is set by the coherence time of the infrared light. For typical laser [57] and plasma parameters [58], the resonant heating of the plasma leads to a drop in the transmitted infrared power that can be readily observed [59]
$\frac{\Delta P_{\rm IR}}{P_{\rm IR}}$ $=$ $(6.5\pm 0.8)\%\left(\frac{\Delta\nu_{\rm IR}}{1~{}{\rm MHz}}\right)^{-1}\left(\frac{\beta}{0.8}\right)$
$\times$ $\left(\frac{T}{12,000~{}{\rm K}}\right)^{-1}\left(\frac{L}{100~{}{\rm cm}}\right)^{2}\left(\frac{B}{0.32~{}{\rm T}}\right)^{-2},$
where, $\Delta\nu_{\rm IR}$ is the coherent line-width of the infrared laser, $\beta$ is the ratio of the electron pressure to the magnetic energy density in the plasma, $T$ is the temperature of the plasma, and $L$ is the laser path length through the plasma.
Microwave Haloscopy—Figure 2 also shows how the infrared magnetic field $B_{\rm IR}$ of laser photons passing through a microwave cavity converts under the influence of axions from the local dark matter halo of the galaxy into the radio-frequency electric field $E_{\rm RF}$ that excites the cavity to resonate with power [56]
$P_{\rm RF}=\frac{1}{2}P_{\rm IR}\left(\frac{\Delta a_{\mu}}{a_{\mu}}\right)^{2}\frac{1}{N_{\mu}}\frac{V_{\rm IR}}{V_{\mu}}\frac{T_{\rm IR}}{T_{\mu}},$ (19)
where, $P_{\rm IR}$ is the infrared laser power, $V_{\rm IR}$ is the volume of the infrared laser beam within the microwave cavity, and $T_{\rm IR}$ is the coherence time of the infrared laser. For a receiver at noise temperature $T_{S}$ and a search that takes total time $T_{A}$, the radio-frequency power gives resonant signal-to-noise ratio [60]
${\rm SNR}$ $=$ $(10.0\pm 1.2)\left(\frac{P_{\rm IR}}{3~{}{\rm mW}}\right)\left(\frac{\Delta\nu_{\rm IR}}{1~{}{\rm MHz}}\right)^{-1}\left(\frac{D}{1~{}{\rm cm}}\right)^{2}$ (20)
$\times$ $\left(\frac{T_{A}}{10^{6}~{}{\rm sec}}\right)^{1/2}\left(\frac{L}{112~{}{\rm cm}}\right)\left(\frac{T_{S}}{77~{}{\rm K}}\right)^{-1},$
where, $\Delta\nu_{\rm IR}=1/(2\pi T_{\rm IR})$ is the infrared laser line-width, $D$ is the beam diameter, and $L$ is the length of the cavity. For typical laser [57] and cavity [61] parameters, the mass of the axion can be determined with precision $h\Delta\nu_{\rm IR}/(m_{A}c^{2})=8.2~{}{\rm ppb}$ that is more than double that of the mass of the muon $\Delta m_{\mu}/m_{\mu}=22~{}{\rm ppb}$ [12] in a search time $T_{A}=10^{6}~{}{\rm sec}\approx 12~{}{\rm days}$ that takes less than a month.
Bend Magnet Haloscopy— Replacing the infrared magnetic field in Figure 2 with the static magnetic field $B_{\rm DC}$ of a dipole bend magnet and swapping the radio-frequency electric field for an infrared electric field $E_{\rm IR}$, we find the process for converting axions from the local dark matter halo of the galaxy into photons with infrared power [62]
$P_{\rm IR}=S_{\rm IR}\left(\frac{B_{\rm DC}^{2}c^{2}}{2Z_{0}}\right)\left(\frac{\Delta a_{\mu}}{a_{\mu}}\right)^{2}\frac{1}{N_{\mu}}\frac{V_{\rm IR}}{V_{\mu}}\frac{T_{\rm IR}}{T_{\mu}},$ (21)
where, $S_{\rm IR}$ is the photosensitive area of the infrared detector, $V_{\rm IR}$ is the volume of the resonant cavity formed by placing infrared mirrors at the entrance and exit of the bend magnet, and $T_{\rm IR}=1/(2\pi\Delta\nu_{\rm IR})$ is the coherence time of the infrared laser with line-width $\Delta\nu_{\rm IR}$ that is used to tune the cavity to the resonance frequency $\omega_{A}=m_{A}c^{2}/\hbar=2\pi\times(122.7\pm 1.0)~{}{\rm THz}$ set by the rest-mass of dark matter. Scanning the infrared frequency through this window of width $\Delta\omega_{A}=2\pi\times 1.0~{}{\rm THz}$ allows the resonance to be found in a search time $T_{A}$ with signal-to-noise ratio [63]
${\rm SNR}$ $=$ $(10.2\pm 1.2)$
$\times$ $\left(\frac{S_{\rm IR}}{2.83\times 10^{-5}~{}{\rm m^{2}}}\right)\left(\frac{B_{\rm DC}}{1~{}{\rm T}}\right)^{2}\left(\frac{D}{1~{}{\rm cm}}\right)^{2}$
$\times$ $\left(\frac{L}{100~{}{\rm cm}}\right)\left(\frac{\Delta\nu_{\rm IR}}{1~{}{\rm MHz}}\right)^{-1/2}\left(\frac{T_{A}}{17.4~{}{\rm hr}}\right)^{1/2}$
$\times$ $\left(\frac{\rm NEP_{A}}{5.3\times 10^{-13}~{}{\rm W/Hz^{1/2}}}\right)^{-1},$
where, $D$ is the laser beam diameter used to tune the cavity, $L$ is the length of the bend magnet, and ${\rm NEP}_{A}$ is the noise equivalent power of the infrared detector at the Compton wavelength of dark matter $\lambda_{A}=(2.44\pm 0.02)~{}{\rm\mu m}$. For typical detector [64] and magnet [65] parameters, the signal-to-noise ratio suffices to directly detect the resonant conversion of dark matter to infrared photons and measure the mass of the axions with precision $2\pi\Delta\nu_{\rm IR}/\omega_{A}=8.2~{}{\rm ppb}$ using a typical line-width $\Delta\nu_{\rm IR}=1~{}{\rm MHz}$ [57] and a search time $T_{A}=17.4~{}{\rm hr}$ taking less than a day.
@@
Phonon Haloscopy— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [66].The incoming mid-infrared light couples to the axion and the coupling creates far-infrared electromagnetic radiation at the resonance frequency $\omega_{\theta}$ of the transverse optical phonon that dominates the far-infrared response of the polar material. The mid-infrared response hits the axion resonance at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ which is detuned by the frequency of the transverse optical phonon $\omega_{\theta}$ from the frequency $\omega_{A}=m_{A}c^{2}/\hbar=2\pi\times(122.7\pm 1.0)~{}{\rm THz}=2\pi c\times(4097\pm 32)~{}{\rm cm^{-1}}$ set by the axion rest-mass.
@@
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ that resonantly excites the polar material provides a table-top determination of the mass of the axion. The size of the dip in mid-infrared transmission through a sample of the polar material on resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient $\alpha(\omega_{\theta})$ [67]:
@@
where $\Gamma_{A\theta}$ is the width of the axion resonance and $\Gamma_{\theta}$ is the width of the phonon resonance. Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance is large enough to be readily observable with samples that show a dip in the far-infrared at frequency $\omega_{\theta}$ from the transverse optical phonon [66]
$\alpha(\omega_{A\theta})$ $=$ $4.0\times\alpha(\omega_{\theta})$ (24)
$\times$ $\left(\frac{\Gamma_{A\theta}}{2000~{}{\rm cm^{-1}}}\right)\left(\frac{\Gamma_{\theta}}{10~{}{\rm cm^{-1}}}\right)^{-1}$
$\times$ $\left(\frac{\omega_{\theta}}{80~{}{\rm cm^{-1}}}\right)\left(\frac{\omega_{A\theta}}{4016~{}{\rm cm^{-1}}}\right)^{-1}.$
@@
Conclusion— We conclude that the tensions between measurements of muon spin precession and standard model calculations are due to axions from the local dark matter halo of the galaxy. The resolution of the muon tension rests on a novel physical picture for the nature of dark matter. The picture leads to percent level predictions for the mass of the axion and for its coupling to electromagnetic fields.
Using kaon decays, we showed that dark matter in the local halo of the galaxy is made of axions. The observed charge asymmetry in kaon decays agrees with the asymmetry created by axions provided the axions have energy density that matches that of the local dark matter halo. The agreement suggests that the kaon decay asymmetry is due to axions in the local dark matter halo and that the local halo is made of axions.
The shift in the muon spin precession frequency due to axions from the local dark matter halo depends not just on the local axion energy density but also on the axion mass and electromagnetic coupling strength. For the mass of the axion predicted by our novel physical picture of the nature of dark matter, the predicted shift in the muon spin precession resolves the tension between precession measurements and Standard Model calculations. In addition, axions from the local dark matter halo convert the neutral vector meson that dominates photon-hadron interactions into its slightly heavier and less dominant isoscalar cousin. The resulting axion-driven shift of the measurements of the leading order hadronic contribution to muon spin precession resolves the $5.1\sigma$ discrepancy between the measurements that dominated the previous consensus [6, 44, 45, 46, 47] and the more recent measurements that are part of the emerging consensus for the Standard Model value [35, 8].
We have proposed table-top experiments to confirm that dark matter is made of axions. We predict a temperature shift of the proton spin precession frequency in liquids and a density shift of the electron transition frequency in optical lattice clocks. The predicted size of both effects is large enough to be seen with presently available table-top experiments.
To directly measure the mass of the axion, we propose to beat dark matter from the local halo against infrared light at the predicted Compton wavelength. The interference creates electromagnetic waves that excite the transverse optical phonon mode in polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ and that drive electron cyclotron motion in a magnetically confined laboratory plasma. The predicted size of the resonant infrared absorption and the predicted size of the electron cyclotron resonance heating in the plasma are both large enough to be seen in currently available table-top experiments.
Finally, we have proposed to use a bend magnet to create infrared photons and to use a microwave cavity to absorb such photons from axions in the local dark matter halo of the galaxy. Measuring the frequency of the infrared photons created by the magnet and absorbed by the cavity gives the mass of dark matter with more than twice the precision of the mass of the muon in searches that take less than a day with the magnet and less than a month with the cavity. These table-top experiments would confirm the proposed form of dark matter with the predicted mass.

### 08 Dec 2023 v36

@@ Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the Standard Model of particle physics. We argue that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument yields a percent level prediction for the mass of the axion provided that dark matter is made of axions. An analysis of charge asymmetry in kaon decays suggests that at least in the local halo dark matter is made of axions and that axions from the local halo are responsible for the observed violation by these reactions of the combined charge conjugation and spatial inversion symmetry operation. A tabletop experiment to directly detect dark matter in the form of axions with the predicted mass is proposed.
Introduction—The spin of the muon precesses in a magnetic field $B$ at frequency $\omega_{a}=a_{\mu}(e/m_{\mu})B$, where, $a_{\mu}\approx\alpha/(2\pi)\approx 1.16\times 10^{-3}$ is the electromagnetic leading order contribution to spin precession, and $e/m_{\mu}=2\pi\times 135~{}{\rm MHz/T}$ is the charge to mass ratio of the muon [1, 2, 3, 4, 5]. In this Letter, we show that axions in the local dark matter halo of the galaxy shift $a_{\mu}$ from the standard model value. The demonstration rests on a novel physical picture for the nature of dark matter which leads to a percent level prediction for the mass of the axion.
@@
Figure 1: Halo axion $A$ collides with virtual photon dressing vertex of positive muon $\mu^{+}$ and magnetic field $B$.
In contrast to the local axion energy density, the axion mass $m_{A}$ and decay constant $f_{A}$ only enter the kaon charge asymmetry through the product $m_{A}f_{A}=\sqrt{\chi_{\rm QCD}}=(5.69\pm 0.05)\times 10^{-3}~{}{\rm GeV^{2}}$ which is fixed by the topological susceptibility $\chi_{\rm QCD}$ of the quantum chromodynamic vacuum [15]. Using the kaon parameters $L_{K}=500~{}{\rm cm}$ and $A_{L}(e)=(3.32\pm 0.08)\times 10^{-3}$ from the current best charge asymmetry measurement [16, 17], the local axion energy density $\rho_{A}=(0.28\pm 0.03)~{}{\rm GeV/cm^{3}}$ that we find from Eq. (3) agrees with the best available estimate $\rho_{\rm DM}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ of the energy density of the local dark matter halo [18]. This agreement suggests that axions saturate the local dark matter halo and that dark matter from the local halo is the source of the charge asymmetry observed in kaon decays.
@@
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [31] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [32]. Tabletop Dark Matter Detector— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [33]. The incoming mid-infrared light couples to the axion and the coupling creates far-infrared electromagnetic radiation at the resonance frequency $\omega_{\theta}$ of the transverse optical phonon that dominates the far-infrared response of the polar material. The mid-infrared response hits the axion resonance at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ which is detuned by the frequency of the transverse optical phonon $\omega_{\theta}$ from the frequency $\omega_{A}=m_{A}c^{2}/\hbar=2\pi\times(122.7\pm 1.0)~{}{\rm THz}=2\pi c\times(4097\pm 32)~{}{\rm cm^{-1}}$ set by the axion rest-mass.
@@
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ that resonantly excites the polar material provides a table-top determination of the mass of the axion. The size of the dip in mid-infrared transmission through a sample of the polar material on resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient [34]:
@@
where $\alpha(\omega_{\theta})$ is the absorption coefficient of the phonon, $\Gamma_{A\theta}\approx 2\pi c/t$ gives the width of the axion resonance in terms of the thickness of the sample $t$, and $\Gamma_{\theta}$ is the width of the phonon resonance. Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance is large enough to be readily observable with samples that show a dip in the far-infrared at frequency $\omega_{\theta}$ from the transverse optical phonon [33]
$\alpha(\omega_{A\theta})$ $=$ $4.0\times\alpha(\omega_{\theta})\left(\frac{t}{5~{}{\rm\mu m}}\right)^{-1}\left(\frac{\Gamma_{\theta}/2\pi c}{10~{}{\rm cm^{-1}}}\right)^{-1}$ (9)
$\times$ $\left(\frac{\omega_{\theta}/2\pi c}{80~{}{\rm cm^{-1}}}\right)\left(\frac{\omega_{A\theta}/2\pi c}{4016~{}{\rm cm^{-1}}}\right)^{-1}.$
@@
Conclusion— We conclude that the tensions between measurements of muon spin precession and standard model calculations are due to axions from the local dark matter halo of the galaxy. The resolution of the muon tension rests on a novel physical picture for the nature of dark matter. The picture leads to percent level predictions for the mass of the axion and for its coupling to electromagnetic fields which can be tested using a tabletop experiment with readily available materials and devices.